# XGBoost-Based Prediction of ICU Mortality in Sepsis-Associated Acute Kidney Injury Patients Using MIMIC-IV Database with Validation from eICU Database

### v1

@@ Abstract
Background: Sepsis-Associated Acute Kidney Injury (SA-AKI) leads to high mortality in intensive care. This study develops machine learning models using the Medical Information Mart for Intensive Care IV (MIMIC-IV) database to predict Intensive Care Unit (ICU) mortality in SA-AKI patients. External validation is conducted using the eICU Collaborative Research Database.
Methods: For 9,474 identified SA-AKI patients in MIMIC-IV, key features like lab results, vital signs, and comorbidities were selected using Variance Inflation Factor (VIF), Recursive Feature Elimination (RFE), and expert input, narrowing to 24 predictive variables. An Extreme Gradient Boosting (XGBoost) model was built for in-hospital mortality prediction, with hyperparameters optimized using GridSearch. Model interpretability was enhanced with SHapley Additive exPlanations (SHAP) and Local Interpretable Model-agnostic Explanations (LIME). External validation was conducted using the eICU database.
Results: The proposed XGBoost model achieved an internal Area Under the Receiver Operating Characteristic curve (AUROC) of 0.878 (95% Confidence Interval: 0.859–0.897). SHAP identified Sequential Organ Failure Assessment (SOFA), serum lactate, and respiratory rate as key mortality predictors. LIME highlighted serum lactate, Acute Physiology and Chronic Health Evaluation II (APACHE II) score, total urine output, and serum calcium as critical features.
Conclusions: The integration of advanced techniques with the XGBoost algorithm yielded a highly accurate and interpretable model for predicting SA-AKI mortality across diverse populations. It supports early identification of high-risk patients, enhancing clinical decision-making in intensive care. Future work needs to focus on enhancing adaptability, versatility, and real-world applications.
@@
Acute Kidney Injury , Sepsis , Machine Learning , XGBoost , Predictive Analytics , External Validation , MIMIC-IV
journal: Computers in Biology and Medicine
@@
[1]organization=Department of Industrial and Systems Engineering, University of Southern California (USC), addressline=3715 McClintock Ave GER 240, city=Los Angeles, postcode=90087, state=California, country=United States
[…]
[2]organization=Colorectal Research Center, Iran University of Medical Sciences, addressline=Tehran Hemat Highway next to Milad Tower, city=Tehran, postcode=14535, state=, country=Iran
[…]
[3]organization=Department of Health Science, California State University, Long Beach (CSULB), addressline=1250 Bellflower Blvd, city=Long Beach, postcode=90840, state=California, country=United States
[…]
[4]organization=Department of Industrial and Manufacturing Engineering, California State Polytechnic University, Pomona, addressline=3801 W Temple Ave, city=Pomona, postcode=91768, state=California, country=United States
{graphicalabstract}
{highlights}
The study implemented a robust machine learning pipeline for predicting ICU mortality in sepsis-associated acute kidney injury (SA-AKI) patients. This pipeline included advanced data preprocessing techniques, stratified imputation for handling missing values, and a three-stage feature selection strategy using Variance Inflation Factor (VIF), Recursive Feature Elimination (RFE), and expert clinical input. The optimized feature set was then used to train an XGBoost model with hyperparameter tuning via GridSearchCV, achieving high predictive accuracy with an AUROC of 0.878 (95% CI: 0.859–0.897) and enhanced clinical applicability. The interpretability analysis using SHAP and LIME identified critical features such as SOFA score, serum lactate, and respiratory rate as key mortality predictors.
The model was externally validated using the eICU Collaborative Research Database, confirming its generalizability and robustness across diverse patient populations with an AUROC of 0.720 (95% CI: 0.708–0.733). This transparent, data-driven approach supports early identification of high-risk patients, optimizing clinical decision-making and resource allocation in intensive care settings.
1 Background
@@
Notable examples of scoring systems used in ICU to assess the severity of illness include the Acute Physiology and Chronic Health Evaluation II (APACHE II) [10], the Simplified Acute Physiology Score II (SAPS II) [11], and the Sequential Organ Failure Assessment (SOFA) score [12]. However, these systems have inherent limitations, particularly in achieving optimal specificity and sensitivity, which may compromise their accuracy in predicting patient outcomes. [13]. Emerging researches highlight the growing integration of machine learning (ML) in clinical predictive analysis, facilitating the identification of complex patterns in multivariate data sets. Advanced ML models such as XGBoost, RandomForest, CatBoost, LightGBM, Support Vector Classifier (SVC), and Logistic Regression have demonstrated substantial effectiveness. Among these methods, XGBoost has gained considerable recognition for its exceptional performance in predictive modeling, particularly in clinical applications. Its capacity to effectively handle large, complex datasets, model intricate feature interactions, and maintain robustness against overfitting has made it a valuable tool for tasks such as disease prediction and patient outcome forecasting. For instance, Ashrafi et al. (2024) employed XGBoost to develop a prediction model for heart failure mortality in ICU patients, achieving a test AUROC of 0.9228, significantly surpassing both prior models and those reported in existing literature [14]. Similarly, Yang et al. (2023) developed and validated a ML-based model using XGBoost to make predictions of 90-day mortality in ICU trauma patients, achieving an accuracy of 0.828, and demonstrating superior calibration and clinical applicability compared to traditional logistic regression [15]. Other machine learning models have also demonstrated high performance in clinical predictive tasks. For example, Miao et al. (2024) applied the LightGBM model to predict ICU readmissions in patients with intracerebral hemorrhage (ICH), achieving an AUROC of 0.736, outperforming other models [16]. Furthermore, Yang et al. (2023) developed a machine learning-based prediction model for Acute Respiratory Distress Syndrome (ARDS) in ICU patients, with CatBoost, known for its effective handling of categorical data and reduced overfitting, emerging as the top performer, achieving an AUROC of 0.817 and demonstrating strong predictive capabilities for ARDS risk stratification [17].
@@
2 Methods
2.1 Data Source
@@ 2.5 Statistic Analysis
Two separate statistical analyses were conducted to assess the variations between the training and test datasets. The first aimed to determine the representativeness of the training dataset relative to the test dataset. The second analysis focused on distinguishing survivors from non-survivors, aiming to identify factors that may be linked to patient outcomes. Two-sided t-tests were employed exclusively to compare group means. Statistical significance was defined as a P-value of less than 0.05. This targeted strategy enabled a comprehensive evaluation of dataset attributes and revealed significant differences that might impact downstream analyses.
2.6 Modeling
@@
The proposed model’s performance was assessed using a train-test split, allocating 75% of the dataset for training and the remaining 25% for testing. This method allows the model to learn from a comprehensive dataset while being evaluated on unseen data, offering an accurate measure of its generalization potential. Additionally, this methodology allows for early detection of potential overfitting, ensuring that the model performs effectively on real-world data.
The study population comprised 83.8% survivors and 16.2% non-survivors, indicating a significant class imbalance in the dataset. To address this issue, a combination of resampling techniques was employed, specifically SMOTE (Synthetic Minority Oversampling Technique) and ADASYN (Adaptive Synthetic Sampling). SMOTE is a widely used method in recent studies, generating synthetic samples by interpolating between existing minority class instances [27, 28, 29]. ADASYN, also popular in contemporary research, adaptively focuses on minority class samples that are more difficult to learn [30, 31], further refining this process by focusing on regions where the minority class is most underrepresented.
@@
The proposed model used in this study was an XGBoost, a machine learning algorithm based on gradient boosting. Known for its efficiency and accuracy, XGBoost is capable of handling both small and large datasets, capturing non-linear relationships, and providing robust predictions. Its ability to manage missing values and support feature importance analysis makes it particularly suitable for clinical data analysis.
@@
Hyperparameter optimization for the model was conducted using a Grid Search approach, which systematically explores various combinations of hyperparameter values to determine the configuration that yields the best performance for the XGBClassifier. This tuning process significantly improved the model’s predictive accuracy and operational efficiency, ensuring optimal performance for the task.
@@
To comprehensively evaluate the performance of the proposed model, five widely utilized machine learning algorithms—LightGBM, SVC, CatBoost, Random Forest, and Logistic Regression—were implemented as baseline models. These algorithms were selected due to their demonstrated efficiency in classification tasks and their capability to manage multivariate and intricate datasets. Each model was trained and tested using the same train-test split to ensure a standardized and fair comparison. Key metrics, including predictive accuracy, precision, sensitivity were calculated for all models to provide a detailed assessment of their performance. This structured evaluation highlights the strengths of the proposed XGBoost model relative to both traditional and advanced machine learning techniques. The results from the baseline models not only establish a benchmark for the proposed approach but also offer critical insights into its comparative effectiveness in handling the dataset’s complexities and achieving reliable classification outcomes.
3 Results
3.1 Statistical Evaluation
@@
T-test analysis [32] comparing the training and test datasets showed that all 24 features had P-values above 0.05, with the smallest value exceeding 0.21, indicating no statistically significant differences between the two datasets. Conversely, the comparison between survivors and non-survivors demonstrated statistically significant differences in 23 out of the 24 features, with P-values approaching 0. The only exception was Creatinine, which exhibited a P-value above 0.05. This comprehensive statistical evaluation provided valuable insights into the dataset characteristics, identifying features associated with survival outcomes while confirming the representativeness of the training and test datasets. Table 2 provides a summary of the population characteristics for the training and test datasets, while the comparison between survivors and non-survivors is displayed in Table 3.
@@
Creatinine 2.580 [1.000–7.850] 2.584 [1.097–6.214] 0.942
3.2 Structure and Performance of Models
@@
3.3 Interpretability
3.3.1 Feature Importance Analysis Using SHAP
@@
3.3.2 Local Feature Contributions to Mortality Risk Using LIME
@@
3.4 Comparison with Best Existing Literature
@@
Table 6: Performance Metrics for Models (Internal Validation)

### v2

# Machine Learning-Based Prediction of ICU Mortality in Sepsis-Associated Acute Kidney Injury Patients Using MIMIC-IV Database with Validation from eICU Database

@@ Abstract
Sepsis-Associated Acute Kidney Injury (SA-AKI) leads to high mortality in intensive care. This study develops machine learning models using the Medical Information Mart for Intensive Care IV (MIMIC-IV) database to predict Intensive Care Unit (ICU) mortality in SA-AKI patients. External validation is conducted using the eICU Collaborative Research Database.
For 9,474 identified SA-AKI patients in MIMIC-IV, key features like lab results, vital signs, and comorbidities were selected using Variance Inflation Factor (VIF), Recursive Feature Elimination (RFE), and expert input, narrowing to 24 predictive variables. An Extreme Gradient Boosting (XGBoost) model was built for in-hospital mortality prediction, with hyperparameters optimized using GridSearch. Model interpretability was enhanced with SHapley Additive exPlanations (SHAP) and Local Interpretable Model-agnostic Explanations (LIME). External validation was conducted using the eICU database.
The proposed XGBoost model achieved an internal Area Under the Receiver Operating Characteristic curve (AUROC) of 0.878 (95% Confidence Interval: 0.859–0.897). SHAP identified Sequential Organ Failure Assessment (SOFA), serum lactate, and respiratory rate as key mortality predictors. LIME highlighted serum lactate, Acute Physiology and Chronic Health Evaluation II (APACHE II) score, total urine output, and serum calcium as critical features.
The integration of advanced techniques with the XGBoost algorithm yielded a highly accurate and interpretable model for predicting SA-AKI mortality across diverse populations. It supports early identification of high-risk patients, enhancing clinical decision-making in intensive care. Future work needs to focus on enhancing adaptability, versatility, and real-world applications.
@@
Introduction
@@
Notable examples of scoring systems used in ICU to assess the severity of illness include the Acute Physiology and Chronic Health Evaluation II (APACHE II) [10], the Simplified Acute Physiology Score II (SAPS II) [11], and the Sequential Organ Failure Assessment (SOFA) score [12]. However, these systems have inherent limitations, particularly in achieving optimal specificity and sensitivity, which may compromise their accuracy in predicting patient outcomes. [13]. Emerging researches highlight the growing integration of machine learning (ML) in clinical predictive analysis, facilitating the identification of complex patterns in multivariate data sets. Advanced ML models such as XGBoost, RandomForest, CatBoost, LightGBM, Support Vector Classifier (SVC), and Logistic Regression have demonstrated substantial effectiveness. Among these methods, XGBoost has gained considerable recognition for its exceptional performance in predictive modeling, particularly in clinical applications. Its capacity to effectively handle large, complex datasets, model intricate feature interactions, and maintain robustness against overfitting has made it a valuable tool for tasks such as disease prediction and patient outcome forecasting. For instance, Ashrafi et al. (2024) employed XGBoost to develop a prediction model for heart failure mortality in ICU patients, achieving a test AUROC of 0.9228, significantly surpassing both prior models and those reported in existing literature [14]. Similarly, Tian et al. (2024) developed an XGBoost-based model to predict AKI in liver cirrhosis patients, achieving an AUROC of 0.832. Their work demonstrates the efficacy of machine learning in renal complication prediction [15]. Other machine learning models have also demonstrated high performance in clinical predictive tasks. For example, Miao et al. (2024) applied the LightGBM model to predict ICU readmissions in patients with intracerebral hemorrhage (ICH), achieving an AUROC of 0.736, outperforming other models [16]. Furthermore, Safaei et al. (2022) developed a machine learning-based prediction model for mortality in ICU patients, with CatBoost, known for its effective handling of categorical data and reduced overfitting, emerging as the top performer, achieving an AUROC of 0.86-0.92 and demonstrating strong predictive capabilities for ARDS risk stratification [17].
@@
Methods
Data Source
@@ Methods
Two separate statistical analyzes were performed to assess the variations between the training and test data sets. The first aimed to determine the representativeness of the training data set relative to the test data set. The second analysis focused on distinguishing survivors from non-survivors, aiming to identify factors that may be linked to patient outcomes. Two-sided t-tests were employed exclusively to compare group means. Statistical significance was defined as a P-value of less than 0.05. This targeted strategy enabled a comprehensive evaluation of dataset attributes and revealed significant differences that might impact downstream analyses.
Modeling
The performance of the proposed model was evaluated using a train-test split, which allowed 75% of the data set to be used for training and the remaining 25% for testing.
@@
This method allows the model to learn from a comprehensive data set while being evaluated on unseen data, offering an accurate measure of its generalization potential. Additionally, this methodology allows for early detection of potential overfitting, ensuring that the model performs effectively on real-world data.
The data set exhibited a significant class imbalance, with only 16.2% instances belonging to the minority class. To mitigate this issue, we employ two widely used oversampling techniques: Synthetic Minority Oversampling Technique (SMOTE) and Adaptive Synthetic Sampling (ADASYN).
SMOTE generates new synthetic samples by interpolating between a given sample of minority class and one of its $k$-nearest neighbors. This method has been widely applied in recent clinical machine learning studies [27, 28, 29]. Mathematically, a synthetic sample $x_{\text{new}}$ is created as:
$x_{\text{new}}=x_{i}+\delta\cdot(x_{nn}-x_{i}),\quad\delta\sim\mathcal{U}(0,1)$ (1)
where $x_{i}$ is a minority class instance, $x_{nn}$ is a randomly selected neighbor from its $k$-nearest neighbors, and $\delta$ is a random value drawn from the uniform distribution $\mathcal{U}(0,1)$.
ADASYN further enhances the oversampling process by adaptively focusing on samples that are more difficult to learn. Adjusts the number of synthetic samples generated for each minority instance based on the local class distribution. This approach has also shown effectiveness in contemporary health informatic studies [30, 31]. The number of synthetic samples $G_{i}$ to generate for a given instance $x_{i}$ is calculated as:
$G_{i}=\frac{r_{i}}{\sum_{j=1}^{n}r_{j}}\cdot G$ (2)
where $r_{i}$ is the estimate of the local density of the difficult samples and $G$ is the total number of synthetic samples to generate.
@@
The proposed model used in this study was an XGBoost, a machine learning algorithm based on gradient boosting. Known for its efficiency and accuracy, XGBoost is capable of handling both small and large datasets, capturing nonlinear relationships, and providing robust predictions. Its ability to manage missing values and support the analysis of features is particularly suitable for clinical data analysis.
@@
Hyperparameter optimization for the model was performed using a grid search approach, which systematically explores various combinations of hyperparameter values to determine the configuration that yields the best performance for the XGBClassifier. This mechanism can be abstracted as searching in the parameter space:
$\text{Grid}=\{(\eta,\text{max\_depth},\text{n\_estimators})\mid\eta\in\mathcal{L}_{1},\text{max\_depth}\in\mathcal{L}_{2},\text{n\_estimators}\in\mathcal{L}_{3}\}$ (3)
This tuning process significantly improved the model’s predictive accuracy and operational efficiency, ensuring optimal performance for the task.
@@
To comprehensively evaluate the performance of the proposed model, five widely utilized machine learning algorithms—LightGBM, SVC, CatBoost, Random Forest, and Logistic Regression—were implemented as baseline models. These algorithms were selected because of their demonstrated efficiency in classification tasks and their ability to manage complex and multivariate datasets. Each model was trained and tested using the same train-test split to ensure a standardized and fair comparison. Key metrics, including predictive accuracy, precision, and sensitivity, were calculated for all models to provide a detailed assessment of their performance. This structured evaluation highlights the strengths of the proposed XGBoost model in relation to both traditional and advanced machine learning techniques. The results of the baseline models not only establish a benchmark for the proposed approach but also offer critical information on its comparative effectiveness in handling the complexities of the dataset and achieving reliable classification results.
Results
Statistical Evaluation
@@ Results
T-test analysis [32] comparing the training and test datasets showed that all 24 features had P-values above 0.05, with the smallest value exceeding 0.21, indicating no statistically significant differences between the two datasets. Conversely, the comparison between survivors and non-survivors demonstrated statistically significant differences in all the 24 features, with P-values approaching zero. The only exception was Creatinine, which exhibited a P-value of 0.002. This comprehensive statistical evaluation provided valuable insights into the dataset characteristics, identifying features associated with survival outcomes while confirming the representativeness of the training and test datasets. Table 2 provides a summary of the population characteristics for the training and test datasets, while the comparison between survivors and non-survivors is displayed in Table 3.
@@
Creatinine 2.580 [1.000–7.850] 2.584 [1.097–6.214] 0.002
Structure and Performance of Models
@@
Interpretability
Feature Importance Analysis Using SHAP
@@
Local Feature Contributions to Mortality Risk Using LIME
@@
Comparison with Best Existing Literature
@@
Table 6: Comparison of Performance Metrics